A straight wire in a magnetic field experiences a force of 0.022 N when the current in the wire is 1.1 A. The current in the wire is changed, and the wire experiences a force of 0.067 N as a result. What is the new current?


1
Expert's answer
2021-03-23T11:09:10-0400

The magnetic force that acts on the wire in the first case can be written as follows:


"F_1=BI_1Lsin\theta."

The magnetic force that acts on the wire in the second case can be written as follows:


"F_2=BI_2Lsin\theta."

Let's divide the first equation by the second one:


"\dfrac{F_1}{F_2}=\dfrac{I_1}{I_2},""I_2=\dfrac{F_2I_1}{F_1}=\dfrac{0.067\ N\cdot1.1\ A}{0.022\ N}=3.35\ A."